# See it. Say it. Sorted: Agentic System for Compositional Diagram Generation

### v1

@@ 1 Introduction
Diffusion-based generative models achieve high-fidelity image synthesis from text and sketches Zhang and Agrawala (2023); Kawar et al. ; Mou et al. (2023), but struggle with the compositional reasoning and spatial precision needed for structured diagrams. Their pixel-based diffusion process makes it difficult to maintain distinct shapes, lines, and alignments Koley et al. (2023); Zhang et al. (2024), and they are not easily adaptable to varying diagram sizes or fine-grained user instructions.
Meanwhile, Vision–Language Models (VLMs) and Large Language Models (LLMs) have shown strong 2D spatial reasoning Yamada et al. (2023); Wu et al. (2024); Cai et al. . Agentic VLM+LLM systems have been applied to visual-feedback tasks like 3D graphics editing Huang et al. (2024); Gu et al. (2025), part assembly Yamada et al. (2024), and scene generation Hu et al. (2024), with the VLM acting as critic and the LLM generating code. Other works train transformers to produce graphics formats like Scalable Vector Graphics (SVG) and Constructive Solid Geometry (CSG) Carlier et al. (2020); Wu et al. ; Kapur et al. (2024), though these lack the generalization of frontier VLMs.
@@
At each optimization step, the Critic VLM compares the target and current images, then:
  1. 1.
    Gives a high-level scene description.
  2. 2.
    Identifies 1–3 key discrepancies.
  3. 3.
    Suggests targeted modifications.
@@
  • VLMs are more reliable at describing qualitative relations than at estimating precise numerical values Chen et al. (2024).
@@ 4 Discussion
We presented See it. Say it. Sorted., a training-free VLM+LLM agentic system that transforms hand-drawn sketches into structured, editable SVG diagrams through an iterative Critic–Candidates–Judge loop. By emphasizing qualitative, relational feedback rather than brittle numerical estimates, the method preserves global structural constraints such as alignment and connectivity while enabling precise local refinements, outperforming frontier image LLMs in faithfully reconstructing compositional layouts.
Because outputs are programmatic, the system is inherently extensible to real-world graphics design environments such as PowerPoint or Google Slides via APIs, and can incorporate additional computer graphics utilities—e.g., auto-layout, snapping, arrow libraries, style transfer—to expand aesthetic control while maintaining semantic correctness. As VLMs and LLMs continue to advance, the same architecture will naturally improve without retraining, offering increasing accuracy and efficiency. In current experiments, the primary bottleneck lies in the Critic VLM, which occasionally produces imprecise or inappropriate suggestions that can lead to worse candidates; this is mitigated by the Judge VLM, which filters regressions and ensures optimization stability. Beyond this application, VLM-based judging Zheng et al. (2023) represents a promising approach for broader agentic systems and reinforcement learning in visual domains, but systematic evaluations of judge reliability—covering calibration, robustness, and domain generalization—are needed. Looking ahead, the qualitative-critique $\rightarrow$ multi-strategy synthesis $\rightarrow$ judging loop underlying our method could extend naturally to 3D tasks such as part assembly, scene layout Liu et al. (2025), and CAD editing Jones et al. (2021), leveraging recent advances in 3D spatial reasoning.

### v2

# Space in Vision, Language, Embodied AI

@@ 1 Introduction
Diffusion-based generative models achieve high-fidelity image synthesis from text and sketches [17, 9, 12], but struggle with the compositional reasoning and spatial precision needed for structured diagrams. Their pixel-based diffusion process makes it difficult to maintain distinct shapes, lines, and alignments [10, 18], and they are not easily adaptable to varying diagram sizes or fine-grained user instructions.
Meanwhile, Vision–Language Models (VLMs) and Large Language Models (LLMs) have shown strong 2D spatial reasoning [15, 14, 1]. Agentic VLM+LLM systems have been applied to visual-feedback tasks like 3D graphics editing [6, 4], part assembly [16], and scene generation [5], with the VLM acting as critic and the LLM generating code. Other works train transformers to produce graphics formats like Scalable Vector Graphics (SVG) and Constructive Solid Geometry (CSG) [2, 13, 8], though these lack the generalization of frontier VLMs.
@@
At each optimization step, the Critic VLM compares the target and current images, then: 1. Gives a high-level scene description; 2. Identifies 1–3 key discrepancies; 3. Suggests targeted modifications.
@@
  • VLMs are more reliable at describing qualitative relations than at estimating precise numerical values [3].
@@ 4 Discussion
We presented See it. Say it. Sorted., a training-free VLM+LLM agentic system that transforms hand-drawn sketches into structured, editable SVG diagrams through an iterative Critic–Candidates–Judge loop. By emphasizing qualitative, relational feedback rather than brittle numerical estimates, the method preserves global structural constraints such as alignment and connectivity while enabling precise local refinements, outperforming frontier image LLMs in faithfully reconstructing compositional layouts. Because outputs are programmatic, the system is inherently extensible to real-world graphics design environments such as PowerPoint or Google Slides via APIs, and can incorporate additional computer graphics utilities—e.g., auto-layout, snapping, arrow libraries, style transfer—to expand aesthetic control while maintaining semantic correctness. As VLMs and LLMs continue to advance, the same architecture will naturally improve without retraining, offering increasing accuracy and efficiency. In current experiments, the primary bottleneck lies in the Critic VLM, which occasionally produces imprecise or inappropriate suggestions that can lead to worse candidates; this is mitigated by the Judge VLM, which filters regressions and ensures optimization stability. Beyond this application, VLM-based judging [19] represents a promising approach for broader agentic systems and reinforcement learning in visual domains, but systematic evaluations of judge reliability—covering calibration, robustness, and domain generalization—are needed. Looking ahead, the qualitative-critique $\rightarrow$ multi-strategy synthesis $\rightarrow$ judging loop underlying our method could extend naturally to 3D tasks such as part assembly, scene layout [11], and CAD editing [7], leveraging recent advances in 3D spatial reasoning.